# Externally Valid Policy Evaluation Combining Trial and Observational Data

### v2

@@ 1 Introduction
Randomized controlled trials (rct) are often considered to be the ‘gold standard’ when evaluating the effects of different decisions or, more generally, decision policies. rct studies circumvent the need to identify and model potential confounding variables that arise in observational studies. Trial studies enable the evaluation and learning of decision policies for use in, e.g., clinical decision support, precision medicine and recommendation systems (Qian & Murphy, 2011; Zhao et al., 2012; Kosorok & Laber, 2019).
However, rcts sample individuals that may differ systematically from a target population of interest. For instance, clinical trials usually involve only individuals who do not have any relevant comorbidities and those who volunteer for trials may very well exhibit different characteristics than the target population. It is also common than studies are limited to a specific geographical area, even when the intended target is a broader region. Invalid inferences about a decision policy can be potentially harmful in safety-critical applications, where the cautionary principle of “above all, do no harm” applies (Smith, 2005). This is especially challenging since the distributions of population characteristics are unknown. How can we generalize results from the trial sample to the intended population?
The focus of this paper is the problem of establishing externally valid inferences about outcomes in a target population, when using experimental results from a trial population (Campbell & Stanley, 1963; Manski, 2007; Westreich, 2019). We consider evaluating a decision policy, denoted $\pi$, that maps covariates $X$ of an individual onto a recommended action $A$. The outcome of this decision has an associated loss $L$ (aka. disutility or negative reward). We assume the availability of samples $(X,A,L)$ from the trial population as well as covariate data $X$ from the target population (Lesko et al., 2016; Li et al., 2022; Colnet et al., 2020). The covariate data is used to model the sampling of individuals in the trial study. We propose a method for evaluating policy $\pi$ that
@@
  • takes into account possible covariate shifts from trial to target distribution, even when using miscalibrated sampling models,
@@
The rest of the paper is outlined as follows. We first state the problem of interest in Section 2 and relate it to the existing literatures in Section 3. We then propose a policy evaluation method in Section 4 and demonstrate its properties using both synthetic and real data in Section 5. We conclude the paper with a discussion in Section 6 about the properties of the method and some future directions of research.
@@ 2 Problem Formulation
Any policy $\pi$, whether deterministic or randomized, can be described by a distribution $p_{\pi}(A|X)$. Each covariate $X\in\mathbb{R}^{p}$ and action $A$ has an associated loss $L\in\mathbb{R}$. We consider here a discrete action space, i.e. $A\in\{1\dots,K\}$. The decision process has a causal structure that can be formalized by a directed acyclic graph, visually summarized in Figure 2 (Peters et al., 2017). The sampling indicator $S$ indicates whether individuals are drawn from a target population, $S=0$, or a trial population, $S=1$. This may depend on their characteristics associated with $X$. The causal structure allows to decompose the two distributions. The target distribution factorizes as
$p_{\pi}(X,A,L|S=0)=p(X|S=0)\cdot p_{\pi}(A|X)\cdot p(L|A,X),$ (1)
@@
$p(X,A,L|S=1)=p(X|S=1)\cdot p(A|X)\cdot p(L|A,X),$ (2)
where only the randomization mechanism $p(A|X)$ is known. In general the characteristics of target and trial populations may differ, that is, $p(X|S=0)\neq p(X|S=1)$.
$X$$A$$L$$S\!\!=\!\!0$
$X$$A$$L$$S\!\!=\!\!1$
Figure 2: Causal structure of process (a) under policy $\pi$ as well as (b) trial study. Sampling indicator $S$ distinguishes between the two. For the important case of rct, assignment of $A$ is not influenced by any covariates so that the path $X\rightarrow A$ is eliminated.
From the trial distribution equation 2, we sample $m$ individuals
$\mathcal{D}=\big{(}(X_{i},A_{i},L_{i})\big{)}^{m}_{i=1}$ (3)
independently. In addition, we also obtain $n$ independent samples of covariate-only data $\big{(}X_{1},X_{2},\dots,X_{n}\big{)}$ from the target population equation 1. Our aim is to infer the out-of-sample loss $L_{n+1}$ for individual $n+1$ under any policy $\pi$. Specifically, we seek a loss limit $\ell_{\alpha}$ as a function of $1-\alpha$, such that $L_{n+1}\leq\ell_{\alpha}$ holds with probability $1-\alpha$ as illustrated by the ‘limit curves’ in Figure 0(a).
The sampling pattern of individuals is described by $p(S|X)$. This distribution is unknown, but we assume that a model $\widehat{p}(S|X)$ is available. This model was fitted using held-out data $\{(X_{j},S_{j})\}$ employing either the conventional logistic model or any state-of-the-art machine learning models (as exemplified below). It can also be obtained from previous studies. There is, however, no guarantee that $\widehat{p}(S|X)$ is calibrated and it may indeed diverge from the unknown sampling pattern. Nevertheless, we want inferences about the out-of-sample loss to be valid also for miscalibrated models. We therefore express the degree of miscalibration in terms of the sampling odds:
$\frac{1}{\Gamma}\;\leq\;\underbrace{\frac{p(S=0|X)}{p(S=1|X)}}_{\text{unknown odds}}\bigg{/}\underbrace{\frac{\widehat{p}(S=0|X)}{\widehat{p}(S=1|X)}}_{\text{nominal odds}}\;\leq\;\Gamma.$ (4)
That is, the nominal sampling odds can diverge by a factor $\Gamma$, where $\Gamma=1$ implies a perfectly calibrated model. This model includes all sources of errors (selection bias, model misspecification, estimation error).
@@
The problem we consider is to construct this externally valid limit $\ell^{\Gamma}_{\alpha}$. This limit allows us to infer the full loss distribution of a future individual with $L_{n+1}$ under policy $\pi$, rather than merely the expected loss $\mathbb{E}_{\pi}[L]$. Specifically, the tail losses are important in healthcare and other safety critical applications where erroneous inferences could be harmful, and a cautious approach when implementing new policies is needed. By increasing the range of $\Gamma$, we certify the validity of the inference under increasingly credible assumptions on $\widehat{p}(S|X)$ (Manski, 2003). As the model credibility increases, however, the informativeness of the inferences decreases. Since the upper bound on the losses, $L_{\max}$, is a trivial and uninformative limit, we may define the informativeness of $\ell^{\Gamma}_{\alpha}$ as
@@
That is, the right limit of a limit curve, which decreases with the degree of miscalibration. Figure 0(a) shows curves that are valid for miscalibration in the range $\Gamma\in[1,2]$, where the informativeness is $95\%$ and $85\%$, respectively. The latter figure means that we can infer an upper bound of the loss for $85\%$ of the target population. The resulting informativeness of a limit curve can therefore guide the specified upper bound of the range of miscalibration in any given problem.
Methods for assessing the calibration of models $\widehat{p}(S|X)$ discussed in, e.g., Murphy & Winkler (1977); Naeini et al. (2015); Widmann et al. (2019) can guide the specified lower bound of the range of miscalibration. The nominal odds in equation 4 can be quantized into several bins and for each bin the unknown odds can be estimated by counting the samples from both the target and trial distributions present. In the case of a well-calibrated model, estimated unknown odds should match the quantized nominal odds for each bin. To assess calibration, this process can be iterated across multiple ranges within the dataset and visualized through a reliability diagram, as exemplified in Figure 3. Further details regarding the experimental setup can be found in Section 5.
Figure 3: Reliability diagram of the estimated odds against the average nominal odds $\widehat{p}(S=0|X)/\widehat{p}(S=1|X)$ obtained from logistic and xgboost models $\widehat{p}(S|X)$ for two different target populations A and B. The diagrams provide lower limits for $\Gamma$.
@@ 3 Background Literature
The problem considered herein is related to the problem of causal inference when combining data from randomized controlled trials and observational studies. Examples of the setting where only covariate data from the observational study is available can be found in Lesko et al. (2016) and Li et al. (2022). For additional examples, refer to the survey by Colnet et al. (2020). Within the broader area of generalizability and transportability, the problem represents the case were the sampling probability depends solely on the covariates $X$, and is independent of the action and the loss (Pearl & Bareinboim, 2014; Lesko et al., 2017; Degtiar & Rose, 2023). The problem is also related to a broader literature of statistical learning under covariate shifts, see for instance Shimodaira (2000); Sugiyama et al. (2007); Quinonero-Candela et al. (2008); Reddi et al. (2015); Chen et al. (2016).
@@
This methodology has been applied in various studies, see for example Cole & Stuart (2010); Stuart et al. (2011); Westreich et al. (2017); Buchanan et al. (2018). It is widely recognized that misspecified logistic models can introduce bias when estimating the weights (Colnet et al., 2020) and more recent works have suggested using flexible models, such as generalized boosted methods (Kern et al., 2016), instead. The counterpart to IPSW, used in off-policy evaluation with observational data, is inverse propensity weighting (IPW). The problem with misspecified logistic models also applies here when estimating the classification probabilities, aka. propensity scores (McCaffrey et al., 2004; Lee et al., 2010). In this case, generalized boosted methods and covariate-balancing methods have shown to be promising alternatives (Setodji et al., 2017; Tu, 2019). The literature on IPSW and IPW is mainly focused on average treatment effect estimation, that is $\mathbb{E}_{\pi_{1}}[L]-\mathbb{E}_{\pi_{0}}[L]$ where $\pi_{1}$ and $\pi_{0}$ denote the ‘treat all’ and ‘treat none’ policies, respectively. By contrast, we want to certify the distributional properties of $L_{n+1}$ for any $\pi$, even under miscalibration of $\widehat{p}(S|X)$.
Conformal prediction is a distribution-free methodology focused on creating covariate-specific prediction regions that are valid for finite-samples (Vovk et al., 2005; Shafer & Vovk, 2008). The methodology was extended by Tibshirani et al. (2019) to also work for known covariate shifts. Jin et al. (2023) combined the marginal sensitivity methodology developed in Tan (2006) with the conformal prediction for covariate shifts to perform sensitivity analysis of treatment effects in the case of unobserved confounding. Our methodology draws upon techniques in conformal prediction, but instead of providing covariate-specific predictions intervals under a policy $\pi$, we are concerned with evaluating any $\pi$ over a target population.
The biased sample selection described by equation 4 was considered in the context of policy learning by Lei et al. (2023). In contrast to that work, our primary focus is on ensuring the validity of inferences regarding out-of-sample losses, even when dealing with finite training data. We achieve this using a sample-splitting technique.
@@ 4 Method
Here we construct a limit $\ell^{\Gamma}_{\alpha}(\mathcal{D})$ on the out-of-sample losses under policy $\pi$ that satisfies equation 5 for any given specified degree of miscalibration $\Gamma$. For this we need to describe the distribution shift from trial to target distribution for all samples, including the $n+1$ sample. We begin by considering the true distribution shift expressed using the ratio
$\frac{p_{\pi}(X,A,L|S=0)}{p(X,A,L|S=1)}.$ (8)
Inserting the factorizations equation 1 and equation 2 into this ratio shows that specifying the distribution shift requires the unknown (conditional) covariate distribution $p(X|S)$. We can, however, bound equation 8 using the model of the sampling pattern, $\widehat{p}(S|X)$, as follows:
$c\cdot\underline{W}^{\Gamma}\leq\frac{p_{\pi}(X,A,L|S=0)}{p(X,A,L|S=1)}\leq c\cdot\overline{W}^{\Gamma},$ (9)

### v3

# Externally Valid Policy Evaluation from Randomized Trials Using Additional Observational Data

@@ 1 Introduction
Randomized controlled trials (rct) are often considered to be the ‘gold standard’ when evaluating the effects of different decisions or, more generally, decision policies. rct studies circumvent the need to identify and model potential confounding variables that arise in observational studies. They enable the evaluation and learning of decision policies for use in, e.g., clinical decision support, precision medicine and recommendation systems (Qian and Murphy, 2011; Zhao et al., 2012; Kosorok and Laber, 2019).
However, rcts sample individuals that may differ systematically from a target population of interest. For instance, clinical trials usually involve only individuals who do not have any relevant comorbidities and those who volunteer for trials may very well exhibit different characteristics than the target population. Invalid inferences about a decision policy can be potentially harmful in safety-critical applications, where the cautionary principle of “above all, do no harm” applies (Smith, 2005). This is especially challenging since the distributions of population characteristics are unknown. How can we generalize results from the trial sample to the intended population?
The focus of this paper is the problem of establishing externally valid inferences about outcomes in a target population, when using experimental results from a trial population (Campbell and Stanley, 1963; Manski, 2007; Westreich, 2019). We consider evaluating a decision policy, denoted $\pi$, that maps covariates $X$ of an individual onto a recommended action $A$. The outcome of this decision has an associated loss $L$ (aka. disutility or negative reward). Thus $L$ may directly represent the outcome of interest. We assume the availability of samples $(X,A,L)$ from the trial population and additional covariate data $X$ from the target population (Lesko et al., 2016; Li et al., 2022; Colnet et al., 2024). The covariate data is used to model the sampling of individuals in the trial study. We propose a method for evaluating policy $\pi$ that
@@
  • takes into account possible covariate shifts from trial to target distribution, even when using miscalibrated sampling models with unmeasured selection factors,
@@
The rest of the paper is outlined as follows. We first state the problem of interest in Section 2 and relate it to the existing literature in Section 3. We then propose a policy evaluation method in Section 4 and demonstrate its properties using both synthetic and real data in Section 5. We conclude the paper with a discussion about the properties of the method in Section 6 and its broader impact in Section 7.
@@ 2 Problem Formulation
Any policy $\pi$, whether deterministic or randomized, can be described by a distribution $p_{\pi}(A|X)$. Each covariate $X$, unmeasured selection factor $U$ and action $A$ has an associated loss $L\in(-\infty,L_{\max})$. We consider here a discrete action space, i.e. $A\in\{1\dots,K\}$. The decision process has a causal structure that can be formalized by a directed acyclic graph, visually summarized in Figure 2 (Peters et al., 2017). The sampling indicator $S$ indicates whether individuals are drawn from a target population, $S=0$, or a trial population, $S=1$. The causal structure allows us to decompose the two distributions. Specifically, the target distribution factorizes as
$p_{\pi}(X,U,A,L|S=0)=p(X,U|S=0)\cdot p_{\pi}(A|X)\cdot p(L|X,U,A),$ (1)
@@
$p(X,U,A,L|S=1)=p(X,U|S=1)\cdot p(A|X)\cdot p(L|X,U,A),$ (2)
where only the randomization mechanism $p(A|X)$ is known. In general the characteristics of target and trial populations may differ, that is, $p(X,U|S=0)\neq p(X,U|S=1)$. The unmeasured $U$ may include self-selection factors that are challenging to record. Note, however, that only factors that also affect the loss $L$ are relevant here.
$S\!\!=\!\!0$$X$$U$$A$$L$
$S\!\!=\!\!1$$X$$U$$A$$L$
Figure 2: Causal structure of process (a) under policy $\pi$ as well as (b) the trial study. Sampling indicator $S$ distinguishes between the two. For the important case of rct, assignment of $A$ is not influenced by any covariates so that the path $X\rightarrow A$ is eliminated.
From the trial distribution equation 2, we sample $m$ individuals $\mathcal{D}=\big{(}(X_{i},A_{i},L_{i})\big{)}^{m}_{i=1}$ independently. In addition, we also obtain $n$ independent samples of covariate-only data $\big{(}X_{1},X_{2},\dots,X_{n}\big{)}$ from the target population equation 1. Our aim is to infer the out-of-sample loss $L_{n+1}$ for individual $n+1$ under any policy $\pi$. Specifically, we seek a loss limit $\ell_{\alpha}$ as a function of $1-\alpha$, such that $L_{n+1}\leq\ell_{\alpha}$ holds with probability $1-\alpha$ as illustrated by the ‘limit curves’ in Figure 1(a).
The sampling pattern of individuals is described by $p(S|X,U)$. This distribution is unknown, but we assume that a model $\widehat{p}(S|X)$ is available. This model was fitted using held-out data $\{(X_{j},S_{j})\}$ employing either the conventional logistic model or any state-of-the-art machine learning models (as exemplified below). It can also be obtained from previous studies. There is, however, no guarantee that $\widehat{p}(S|X)$ is calibrated and it may indeed diverge from the unknown sampling pattern. Nevertheless, we want inferences about the out-of-sample loss to be valid also for miscalibrated models. We therefore express the degree of miscalibration in terms of the selection odds:
$\frac{1}{\Gamma}\;\leq\;\underbrace{\frac{p(S=0|X,U)}{p(S=1|X,U)}}_{\text{unknown selection odds}}\bigg{/}\underbrace{\frac{\widehat{p}(S=0|X)}{\widehat{p}(S=1|X)}}_{\text{nominal selection odds}}\;\leq\;\Gamma\qquad\text{(}a.s.)$ (3)
That is, the nominal selection odds can diverge by a factor $\Gamma$, where $\Gamma=1$ implies a perfectly calibrated model. This model includes all sources of errors (selection bias, model misspecification, estimation error).
@@
The problem we consider is to construct this externally valid limit $\ell^{\Gamma}_{\alpha}$. This limit allows us to infer the full loss distribution of a future individual with $L_{n+1}$ under policy $\pi$, rather than merely the expected loss $\mathbb{E}_{\pi}[L]$. Specifically, the tail losses are important in healthcare and other safety critical applications where erroneous inferences could be harmful, and a cautious approach when implementing new policies is needed.
The limit curve $\ell^{\Gamma}_{\alpha}$ for policy $\pi$ is valid up to any declared degree of odds miscalibration $\Gamma$, which establishes the credibility of the policy evaluation, cf. Manski (2003). While $\Gamma$ is unknowable, especially under unmeasured $U$, we can use ideas from sensitivity analysis to guide its selection using measured data (Rosenbaum and Rubin, 1983; Tan, 2006) . Following the method in Huang (2024), we treat observed selection factors in $X$ as unmeasured $U$ in equation 3 to benchmark appropriate values for $\Gamma$, as detailed in subsection 4.1.
By increasing the range of $\Gamma$, we certify the validity of the inference under increasingly credible assumptions on $\widehat{p}(S|X)$ (Manski, 2003). As the model credibility increases, however, the informativeness of the inferences decreases. Since the upper bound on the losses, $L_{\max}$, is a trivial and uninformative limit, we may define the informativeness of $\ell^{\Gamma}_{\alpha}$ as
@@
That is, the right limit of a limit curve, which decreases with the degree of miscalibration. Figure 1(a) shows curves that are valid for miscalibration in the range $\Gamma\in[1,2]$, where the informativeness is $95\%$ and $90\%$, respectively. The latter figure means that we can infer a non-trivial bound on the loss for $90\%$ of the target population.
Remark: This paper addresses the evaluation of a given policy, whether proposed by a clinical expert or learned from historical data. By setting aside samples from an rct study, a policy $\pi$ can be learned, and its out-of-sample performance evaluated using the proposed methodology.
Figure 3: (a) Omitting measured selection factors to benchmark credible values for $\Gamma$ in equation 3. (b) Inferred blood mercury levels [$\mu$g/L] in a target population under ‘high’ and ‘low’ seafood consumption ($\pi_{1}$ and $\pi_{0}$, respectively). Limit curves for degrees of odds miscalibration $\Gamma\in[1,2]$.
@@ 3 Background Literature
The problem considered herein is related to the problem of causal inference when combining data from randomized controlled trials and observational studies. Examples of the setting where only covariate data from the observational study is available can be found in Lesko et al. (2016) and Li et al. (2022). For additional examples, refer to the survey by Colnet et al. (2024). Within the broader area of generalizability and transportability, the problem represents the case were the sampling probability depends solely on the covariates $X$, and is independent of the action and the loss (Pearl and Bareinboim, 2014; Lesko et al., 2017; Degtiar and Rose, 2023). The problem is also related to a broader literature of statistical learning under covariate shifts, see for instance Shimodaira (2000); Sugiyama et al. (2007); Quinonero-Candela et al. (2008); Reddi et al. (2015); Chen et al. (2016).
@@
This methodology has been applied in various studies, see for example Cole and Stuart (2010); Stuart et al. (2011); Westreich et al. (2017); Buchanan et al. (2018). It is widely recognized that misspecified logistic models can introduce bias when estimating the weights (Colnet et al., 2024) and more recent works have suggested using flexible models, such as generalized boosted methods (Kern et al., 2016), instead. The counterpart to IPSW, used in off-policy evaluation with observational data, is inverse propensity weighting (IPW). The problem with misspecified logistic models also applies here when estimating the classification probabilities, aka. propensity scores (McCaffrey et al., 2004; Lee et al., 2010). In this case, generalized boosted methods and covariate-balancing methods have shown to be promising alternatives (Setodji et al., 2017; Tu, 2019). The literature on IPSW and IPW is mainly focused on average treatment effect estimation, that is $\mathbb{E}_{\pi_{1}}[L]-\mathbb{E}_{\pi_{0}}[L]$ where $\pi_{1}$ and $\pi_{0}$ denote the ‘treat all’ and ‘treat none’ policies, respectively. By contrast, we want to certify the distributional properties of $L_{n+1}$ for any $\pi$, even under miscalibration of $\widehat{p}(S|X)$.
Conformal prediction is a distribution-free methodology focused on creating covariate-specific prediction regions that are valid for finite-samples (Vovk et al., 2005; Shafer and Vovk, 2008). The methodology was extended by Tibshirani et al. (2019) to also work for known covariate shifts. Jin et al. (2023) combined the marginal sensitivity methodology developed in Tan (2006) with the conformal prediction for covariate shifts to perform sensitivity analysis of treatment effects in the case of unobserved confounding. Our methodology draws upon techniques in conformal prediction, but instead of providing covariate-specific prediction intervals under a policy $\pi$, we are concerned with evaluating any $\pi$ over a target population.
When full identification of the causal effect is not possible due to unmeasured confounders, partial identification sensitivity analysis can be used to evaluate the robustness of the estimates. Rosenbaum and Rubin (1983) introduced a sensitivity parameter to bound the odds ratio of the probability of treatment. More recent work has extended this approach to account for treatment effect heterogeneity and non-binary treatments, as seen in Tan (2006); Shen et al. (2011); Zhao et al. (2019); Dorn and Guo (2023) among others. However, it is often challenging to interpret the absolute value of the sensitivity parameter in practical scenarios. To address this, recent research has proposed benchmarking, or calibrating, results by estimating the effects of unmeasured confounders, see for example Hsu and Small (2013); Franks et al. (2019); Veitch and Zaveri (2020); De Bartolomeis et al. (2024). Note that all these papers work with sensitivity analysis for observational studies. Huang (2024) instead combines sensitivity analysis for generalization with benchmarking to determine reasonable values of $\Gamma$.
A biased sample selection, similar to that described by equation 3, was considered in the context of average treatment effect estimation by Nie et al. (2021). In contrast to that work, our primary focus is on ensuring the validity of inferences regarding out-of-sample losses, even when dealing with finite training data. We achieve this using a sample-splitting technique.
@@ 4 Method
Here we construct a limit $\ell^{\Gamma}_{\alpha}(\mathcal{D})$ on the out-of-sample losses under policy $\pi$ that satisfies equation 4 for any given specified degree of miscalibration $\Gamma$.
4.1 Benchmarking degree of miscalibration
The limit curve $\ell^{\Gamma}_{\alpha}(\mathcal{D})$ holds up to the specified degree of odds miscalibration $\Gamma$. Although $\Gamma$ is inherently unknown, sensitivity analysis and methods for assessing the calibration of models can guide its estimation using available data.
We start with a benchmarking method that specifically accounts for the potential impact of unmeasured selection factors, $U$. Building on the approach in Huang (2024), we treat observed selection factors in $X$ as proxies for unmeasured $U$ in equation equation 3, providing a benchmark for selecting suitable $\Gamma$ values. Specifically, let the the omitted selection factor $X_{k}$ be $U$ and $X_{-k}$ denotes the remaining covariates. Then the ratio in equation 3 is estimated by dividing $\widehat{\text{odds}}(X_{-k},X_{k})=\frac{\widehat{p}(S=0|X)}{\widehat{p}(S=1|X)}$ by $\widehat{\text{odds}}(X_{-k})=\frac{\widehat{p}(S=0|X_{-k})}{\widehat{p}(S=1|X_{-k})}$. Figure 3(a) summarizes the distribution of $\widehat{\text{odds}}(X_{-k},X_{k})/\widehat{\text{odds}}(X_{-k})$ using a real data set, where the omitted $X_{k}$ is either age, income or education. We see that the corresponding ratios all fall within $\Gamma=2$. We can therefore conclude that if any unmeasured selection factor $U$ is weaker than the omitted factors, it is credible to set $\Gamma$ in the range of $1.5$ to $2$. The corresponding loss curves are shown in Figure 3(b), which illustrates the blood mercury levels in a target population under policies of ‘high’ respective ‘low’ seafood consumption that can credibly be extrapolated from trial data. More details are available in Section 5.2. Note that this benchmarking method serves as a guide for assessing the influence of unmeasured selection factors $U$ (Cinelli and Hazlett, 2019).
Without unmeasured selection factors, methods for assessing the calibration of models $\widehat{p}(S|X)$ discussed in, e.g., Murphy and Winkler (1977); Naeini et al. (2015); Widmann et al. (2019) can guide the specified lower bound of the range of miscalibration. The nominal selection odds in equation 3, i.e., $\widehat{\text{odds}}(X)=\frac{\widehat{p}(S=0|X)}{\widehat{p}(S=1|X)}$, can be quantized into several bins and for each bin the unknown selection odds, i.e., $\text{odds}(X)=\frac{p(S=0|X)}{p(S=1|X)}$, can be estimated by counting the samples from both the target and trial distributions present. In the case of a well-calibrated model, estimated unknown odds should match the quantized nominal odds for each bin. To assess calibration, this process can be iterated across multiple ranges within the dataset and visualized through a reliability diagram, as exemplified in Figure 4. Using equation 3, we have that
Take the expectation with respect to $X$, conditional on the nominal odds in a specified interval (or bin) $I$, so that:
The expected odds is then estimated for each bin $I$ by counting samples from the target and trial distributions.
4.2 Inferring out-of-sample limit
We will now construct the limit $\ell^{\Gamma}_{\alpha}(\mathcal{D})$. For this we need to describe the distribution shift from trial to target distribution for all samples, including the $n+1$ sample. We begin by considering the true distribution shift expressed using the ratio
$\frac{p_{\pi}(X,U,A,L|S=0)}{p(X,U,A,L|S=1)}.$ (9)
Inserting the factorizations equation 1 and equation 2 into this ratio shows that specifying the distribution shift requires the unknown (conditional) covariate distribution $p(X,U|S)$. We can, however, bound equation 9 using the model of the sampling pattern, $\widehat{p}(S|X)$, as follows:
$c\cdot\underline{W}^{\Gamma}\leq\frac{p_{\pi}(X,U,A,L|S=0)}{p(X,U,A,L|S=1)}\leq c\cdot\overline{W}^{\Gamma},$ (10)